ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

20. Special Activities of the ICM 1998

Round Table

1991 MS Classification: 01, 00 Srinivasan, Bhama, University of Illinois at Chicago, U. S. ; Christine Bessenrodt, University of Madgeburg, Germany; Bettye Anne Case, Florida State University, U. S., organizers and moderators Events and policies: Effects on women in mathematics The panelists are women in mathematics from several different countries. Each will discuss impacts she has noted on the work and lives of women in the mathematical sciences which may result from various national policies, practices, and events. Do some of these events cause more women or fewer to participate in mathematics? This panel was organized by a committee representing the Association for Women in Mathematics (AWM), the European Women in Mathematics (EWM), and the Committee on Women and Mathematics of the European Mathematical Society, with further assistance from mathematicians in several additional countries.